# Analyzing Uncertainty Quantification in Statistical and Deep Learning Models for Probabilistic Electricity Price Forecasting

### v1

@@ 4.5 Error Analysis
Next, we examine the PICP metric across different confidence levels to assess how well the predicted intervals are calibrated. The corresponding results are presented in Table 2 and Fig. 4. The naive models consistently produce PICP values higher than the expected nominal coverage, indicating overconfident intervals. The same holds for LEAR-GARCH and LEAR-CP. In contrast, all other models tend to produce underconfident PIs. Both the ensemble and MC Dropout models show improved PICP values compared to the DDNN, especially as the confidence level increases. The greatest improvements are observed around the 9090% confidence level. Up to the 6060% confidence level, the performance of the ensemble and MC Dropout models is quite similar. Moreover, both the ensemble and MC Dropout models significantly reduce the standard deviation of results across independent runs by about half compared to the DDNN, indicating more stable performance. After the 6060% confidence level, the NN-CP models surpass both the ensemble and MC Dropout models. At about the 8080% confidence level, the LEAR-QRA begins to surpass the previous models, while the LEAR-CP achieves the best overall coverage, as reflected in the MAACE metric.
@@
As shown in Table 3 and Figure 5, the MPIW increases almost linearly with the confidence level up to around the 7070% confidence interval. As expected, the naive model produces significantly wider PIs on average, consistent with its high PICP values. In contrast, the other models generate considerably narrower intervals, roughly half the width of those from the naive model. The exception is the EvDNN model, which produces very sharp but less reliable PIs. Notably, LEAR-QRA consistently produces the narrowest PIs across all confidence levels among the traditional and NN-based models (excluding the EvDNN), with MPIW values of $17.88$ at 5050% confidence and about $39.18$ at 8080%. At 9090% and 9898% confidence, however, LEAR-CP and LEAR-GARCH perform best. These values are substantially lower than those of DDNN, which produces intervals of $24.10$ and $58.76$ at the 5050% and 8080% levels, respectively. The ensemble with 5 models shows a high standard deviation across runs, suggesting that its performance is sensitive to the specific composition of the ensemble. Up to the 5050% confidence level, the MPIW values of the ensemble and MC Dropout models (excluding Ens5), as well as the DDNN, are quite similar. Beyond this point, the ensemble and MC Dropout models begin to diverge from the DDNN. At the 9090% confidence level, where the highest gains in PICP are observed, the Ens10 and MC30 models improve PICP by $4.36$ and $5.89$ percentage points, respectively, in exchange for increases in MPIW by $4.14$ and $2.63$, reflecting the trade-off between interval width and coverage. Between the two, MC Dropout performs better overall.
@@ 4.6 Empirical Analysis of Diebold-Mariano Test Results
The empirical p-values are shown as a heatmap in Fig. 6, where each entry $p_{ij}$ denotes the p-value from the DM test comparing model $M_{i}$ (row model) against model $M_{j}$ (column model). The diagonal elements are marked as nan since models are not tested against themselves. The p-values in the DM test show whether the two models under consideration are significantly different in terms of prediction accuracy. However, they do not provide any information about model superiority. To extract this information, one has to analyze the test statistic. The mathematical form of the DM test and the table of the test statistic are given in Appendix .1 in Table 4. Table 4 reports the corresponding DM statistics $d_{i,j}$ (based on CRPS loss), where $d_{i,j}<0$ means $M_{i}$ attains lower average loss than $M_{j}$ (row is better), $d_{i,j}>0$ means the opposite (column is better), $|d_{i,j}|$ quantifies the strength of the advantage, and values numerically near 0 indicate practical equivalence. In Fig. 6, we show the pairwise p-values (row model $M_{i}$ vs. column model $M_{j}$), with significance level $0.05$. Entries whose numeric values on the color bar are <0.05<0.05 indicate that the observed difference between the two models is statistically significant, whereas larger p-values >0.05>0.05 indicate no detectable difference.
Reading both Fig. 6 and Table 4 together, the LEAR versions LEAR-CP and LEAR-GARCH have $d_{i,j}<0$ against essentially all competitors and, for those same pairs, the figure shows color–bar values well below $0.05$ (often close to $0.00$), marking them as the best overall performers. LEAR-QRA is statistically comparable, with predominantly negative $d$-entries and many within-LEAR comparisons showing color–bar values $\geq 0.05$, indicating no detectable differences among the leaders. Next comes EvDNN (mostly negative $d$ against mid-tier models with many $p$-values <0.05<0.05 versus baselines), followed by MCD30 and MCD30-CP, then MCD10. ENS10-CP and Ens10 form a weaker tier with mixed signs and smaller magnitudes (and several comparisons with $p\geq 0.05$). DDNN-CP, DDNN, and ENS5 lag with many positive $d$ against stronger models. Finally, Naive-HStrain and Naive-HSval are decisively the worst, exhibiting large positive $d$ almost everywhere and, correspondingly, numerous $p$-values <0.05<0.05 against virtually all opponents. With 14 models, there are 91 tests; unless otherwise stated, ”significant” here refers to unadjusted $p<0.05$.
In summary, the three LEAR versions are jointly the best: each shows predominantly negative $d$-entries against every non-LEAR model and, for those same pairs, $p<0.05$; among the LEAR models themselves the $d$-entries are near zero and the corresponding $p$-values are $\geq 0.05$, so they are statistically indistinguishable. The next tier is EvDNN, followed by MCD30 and MCD30-CP, then MCD10. Ens10-CP and Ens10 are weaker (mixed signs, smaller $|d|$, many $p\geq 0.05$). DDNN-CP, DDNN, and Ens5 trail, and the naive baselines (Naive-HStrain, Naive-HSval) are decisively the worst, with large positive $d$ almost everywhere and $p<0.05$ against virtually all opponents.

### v2

@@ 4.5 Error Analysis
Next, we examine the PICP metric across different confidence levels to assess how well the predicted intervals are calibrated. The corresponding results are presented in Table 2 and Fig. 4. It should be noted that, of all the figures, there are 10 models under consideration, whereas the tables in this section include 15 models in total. Five models, namely Naive-HStrain, Ens5, MCD10, EvDNN, and EvDNN-CP, are not shown in Fig. 4 and Fig. 5, since we consider only one better-performing model from each category, and the EvDNN model struggles to capture uncertainty. The naive models consistently produce PICP values higher than the expected nominal coverage, indicating underconfident intervals. The same holds for LEAR-GARCH and LEAR-CP. In contrast, all other models tend to produce overconfident PIs. Both the ensemble and MC Dropout models show improved PICP values compared to the DDNN, especially as the confidence level increases. The greatest improvements are observed around the 9090% confidence level. Up to the 6060% confidence level, the performance of the ensemble and MC Dropout models is quite similar. Moreover, both the ensemble and MC Dropout models significantly reduce the standard deviation of results across independent runs by about half compared to the DDNN, indicating more stable performance. After the 6060% confidence level, the NN-CP models surpass both the ensemble and MC Dropout models. At about the 8080% confidence level, the LEAR-QRA begins to surpass the previous models, while the LEAR-CP achieves the best overall coverage, as reflected in the MAACE metric.
@@
As shown in Table 3 and Figure 5, the MPIW increases almost linearly with the confidence level up to around the 7070% confidence interval. As expected, the naive model produces significantly wider PIs on average, consistent with its high PICP values. In contrast, the other models generate considerably narrower intervals, roughly half the width of those from the naive model. The exception is the EvDNN model, which produces very sharp but less reliable PIs. Notably, LEAR-QRA consistently produces the narrowest PIs across all confidence levels among the traditional and NN-based models (excluding the EvDNN), with MPIW values of $17.88$ at 5050% confidence and about $39.18$ at 8080%. At 9090% and 9898% confidence, however, LEAR-CP and LEAR-GARCH perform best. These values are substantially lower than those of DDNN, which produces intervals of $24.10$ and $45.78$ at the 5050% and 8080% levels, respectively. The ensemble with 5 models shows a high standard deviation across runs, suggesting that its performance is sensitive to the specific composition of the ensemble. Up to the 5050% confidence level, the MPIW values of the ensemble and MC Dropout models (excluding Ens5), as well as the DDNN, are quite similar. Beyond this point, the ensemble and MC Dropout models begin to diverge from the DDNN. At the 9090% confidence level, where the highest gains in PICP are observed, the Ens10 and MC30 models improve PICP by $4.36$ and $5.89$ percentage points, respectively, in exchange for increases in MPIW by $4.14$ and $2.63$, reflecting the trade-off between interval width and coverage. Between the two, MC Dropout performs better overall.
@@ 4.6 Empirical Analysis of Diebold-Mariano Test Results
The empirical p-values are shown as a heatmap in Fig. 6, where each entry $p_{ij}$ denotes the p-value from the DM test comparing model $M_{i}$ (row model) against model $M_{j}$ (column model). In this section and the following one, we have not used five models for the Diebold-Mariano (DM) test evaluation and in the trading strategies, namely, Naive-HStrain, Ens5, MCD10, EvDNN, and EvDNN-CP, since we used only one better-performing model from each category, and the EvDNN model struggled to capture uncertainty. The diagonal elements are marked as nan since models are not tested against themselves. The p-values in the DM test show whether the two models under consideration are significantly different in terms of prediction accuracy. However, they do not provide any information about model superiority. To extract this information, one has to analyze the test statistic. The mathematical form of the DM test and the table of the test statistic are given in Appendix .1 in Table 5. Table 5 reports the corresponding DM statistics $d_{i,j}$ (based on CRPS loss), where $d_{i,j}<0$ means $M_{i}$ attains lower average loss than $M_{j}$ (row is better), $d_{i,j}>0$ means the opposite (column is better), $|d_{i,j}|$ quantifies the strength of the advantage, and values numerically near 0 indicate practical equivalence. In Fig. 6, we show the pairwise p-values (row model $M_{i}$ vs. column model $M_{j}$), with significance level $0.05$. Entries whose numeric values on the color bar are <0.05<0.05 indicate that the observed difference between the two models is statistically significant, whereas larger p-values >0.05>0.05 indicate no detectable difference.
In Figure 6 we show the P-values for the DM-test computed at $\alpha=0.05$. It shows that three LEAR–type models form a statistically tied top tier: LEAR–GARCH, LEAR–QRA, and LEAR–CP are mutually indistinguishable, with near–zero differentials and non–rejections in all pairwise directions. The pair wise dm-statistic given in the Table 5 in Appendix .2 supports this claim. For further clarity on the P-values, we also give the numerical values of the P-values in Table 4 (e.g., ${DM}({\text{LEAR-GARCH},\text{LEAR-QRA}})=-0.072$ with $p=0.471$, ${DM}({\text{LEAR-CP},\text{LEAR-GARCH}})=0.049$ with $p=0.520$, and ${DM}({\text{LEAR-CP},\text{LEAR-QRA}})=-0.039$ with $p=0.484$; see Tables 54). $DM(.,.)$ represents the statistic for two models into consideration and is computed as shown in the Appdendix .1 Each of these three models, however, dominates every non-LEAR alternative by a wide and statistically decisive margin: for instance, ${DM}({\text{LEAR-CP},\text{Ens10}})=-4.921$ with $p=0.000$, ${DM}({\text{LEAR-GARCH},\text{DDNN}})=-5.763$ with $p=0.000$, and ${DM}({\text{LEAR-QRA},\text{MCD30-CP}})=-4.839$ with $p=0.000$, and analogous results hold across all such pairs. All non-LEAR models (DDNN, Ens10, MCD30) and their conformal/MC-dropout variants (DDNN-CP, Ens10-CP, MCD30-CP) are statistically indistinguishable from one another at $\alpha=0.05$, however, the directional tendencies do exist (e.g., ${DM}({\text{MCD30},\text{Ens10}})=-1.748$; $({DM}{\text{Ens10},\text{DDNN}})=-2.432$), but the associated pp–values are too large to claim superiority ($p=0.130$ and $p=0.192$, respectively). Conformal models and the MC-dropout models does not yield measurable gains (in terms of CRPS loss) over the corresponding base models: ${DM}({\text{MCD30-CP},\text{MCD30}})=0.438$ with $p=0.624$, ${DM}({\text{Ens10-CP},\text{Ens10}})=0.299$ with $p=0.578$, and ${DM}({\text{DDNN-CP},\text{DDNN}})=-0.495$ with $p=0.422$ all fail to reject. Finally, every method strictly improves upon the naive baseline (e.g., ${DM}({\text{DDNN},\text{Naive-HS}_{val}})=-6.273$, ${DM}({\text{Ens10},\text{Naive-HS}_{val}})=-6.984$, ${DM}(\text{LEAR-CP}$, $\text{Naive-HS}_{val})=-10.235$, all with $p=0.000$). In summary there is no single best model within the LEAR family, but they constitute a statistically superior frontier; (ii) non-LEAR methods are statistically tied with one another and uniformly inferior to LEAR; and (iii) naive forecasting is decisively underperforming.